OpenStudy (anonymous):

The force exerted by a bow string on a 80 gram arrow when it’s fired from the bow is 80N. The string pushes the arrow for .04 seconds. What is the magnitude of the velocity of the arrow when it shots from the bow?

OpenStudy (mathmate):

F=ma, where F=80N, m=80g =0.08kg So a=F/m=1000 m/s Initial velocity: vi=0 Final velocity, vf = vi + at and t= 0.04 s Find vf.

OpenStudy (anonymous):

.04*1000, so 40=final velocity?

OpenStudy (mathmate):

Correct. Are you familiar with the equations shown?

OpenStudy (anonymous):

Yes and thank you!

OpenStudy (mathmate):

You're welcome! :)